a ball rolls off  the edge of  a horizontal table top 2m high. if it strikes the flor at apoint 2 m horizontally away from the edge of the table , what was its speed at the instant it left the table?

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
Vertical displacement = uverticalt+at2/2
=>2=gt2/2
=>t=2/√10 s
Horizontal Displacement= Velocityx time
=>2=vx2/√10
=>v=√10 m/s
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)